On 1 January 2013, Mary Inc issues $1,000,000 face value, 10-year bonds with annual interest rate of 5% to be paid each 31 December. The market interest rate is 4%.

Using the effective interest rate method of amortization, Mary Inc should record when it closes its annual book on Dec 31, 2015

Group of answer choices

a carrying amount of 1,081,109 on Dec 31, 2015

an interest expense of 42,974 on Dec 31, 2015

a cash disbursement of 43,244 on Dec 31, 2015

a carrying amount of 1,060,021 on Dec 31, 2015

an amortization of discount of 7,307 on Dec 31, 2015
Answer #1
A carrying amount of 1,060,021 on Dec 31, 2015
Option D is correct
Workings:
Amount PV factor 4% Present value
Cash Annual interest 50000 8.11090 405545
Principal 1000000 0.67556 675564
Total 1081109
Cash annual interest Interest expense 4% Premium amortized Carrying value
January 01, 2013 1081109
December 31, 2013 50000 43244 6756 1074353
December 31, 2014 50000 42974 7026 1067327
December 31, 2015 50000 42693 7307 1060021
Interest expense 4%:
December 31, 2013 43244 =1081109*4%
December 31, 2014 42974 =1074353*4%
December 31, 2015 42693 =1067327*4%